sin30=

sin30=

Grade:

2 Answers

Kurma Anudeep
11 years ago

sin30=sin45-sin15=1/2

Ramya Ananth
11 years ago

Draw an equilateral triangle ABC.


Draw a perpendicular from A to BC intersecting BC at D
AB = AC (sides of an equilateral trianlge are equal)
Angle ADB = Angle ADC = 90
AD = AD
ΔADB ≅ ΔADC (RHS postulate)

BD = DC = (1/2)BC = (1/2)AB (congruence of triangles and sides of an equilateral triangle are equal)

Angle ABD = AngleDAC = 60/2 = 30 (congruence of triangles)

sinBAD = BD/AB = 1/2
sin(30) = 1/2